OpenStudy (anonymous):

5sqrt y-1-sqrt24y-1=0

OpenStudy (anonymous):

\[5\sqrt{y-1-\sqrt{24y-1}}=0\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sqrt[5]{y-1-\sqrt{24y-1}}=0\] \[y- 1- \sqrt{24y -1} = 0\] \[ \sqrt{24y -1} = 1-y\] \[ {24y -1} = 1-2y+ y^2\]

OpenStudy (anonymous):

the thing wasnt supposed to be so long on the first one.. thats why i tried to draw it lol........

OpenStudy (anonymous):

\[5\sqrt{y-1}-\sqrt{24y-1} = 0\]

OpenStudy (anonymous):

yes thats what it was supposed to be.. im trying to figure out how to use the equation editor.. not going so well

OpenStudy (anonymous):

\[5\sqrt{y-1} = \sqrt{24y-1}\] \[25 \times (y-1) = 24y-1\] \[25y -25 = 24y -1\] \[25y -24y = 25-1\] \[y = 24\]